OpenStudy (anonymous):

How to solve this linear equations -2x+y=0 2x+4y=-8

OpenStudy (anonymous):

y=2x so, sub y+4y=-8 5y=-8 y=-8/5. so, x=-4/5

OpenStudy (precal):